OpenStudy (anonymous):

Give the 10th term in the geometric sequence. 2, 6, 18, 54, ...

OpenStudy (experimentx):

2*3^9

OpenStudy (anonymous):

a = 2 r = 3 a(n) = a*r^(n-1) a(10) = 2*3^(9) ------- a(10) = 39,366